# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 5. Property (U), weak eigenvalue gaps and the uniform gap summation property
$\log\frac{\ell_{i}(\rho^{ss}(\gamma))}{\ell_{i+1}(\rho^{ss}(\gamma))}\geqslant\varepsilon|\gamma|_{\infty}\geqslant\varepsilon c_{2}\big{|}\big{|}\lambda(\rho_{2}(\gamma))\big{|}\big{|}$
for every $\gamma\in\Gamma$. By Theorem 2.5 there exists a finite subset $F$ of $\Gamma$ and $C>0$ such that for every $\gamma\in\Gamma$ there exists $f\in F$ with $\big{|}\big{|}\mu(\rho^{ss}(\gamma))-\lambda(\rho(\gamma f))\big{|}\big{|}\leqslant C$ and $\big{|}\big{|}\mu(\rho_{2}(\gamma))-\lambda(\rho_{2}(\gamma f))\big{|}\big{|}\leqslant C$. In particular, we may choose $R>0$ such that
$\log\frac{\sigma_{i}(\rho^{ss}(\gamma))}{\sigma_{i+1}(\rho^{ss}(\gamma))}\geqslant\varepsilon c_{2}\big{|}\big{|}\mu(\rho_{2}(g))\big{|}\big{|}-R$
for every $g\in\Gamma$. By assumption, for all but finitely many $g\in\Gamma$ we have $\big{|}\big{|}\mu(\rho_{2}(g))\big{|}\big{|}\geqslant 2/(\varepsilon c_{2})\log|g|_{\Gamma}$, so there exists $R^{\prime}>0$ such that
$\log\frac{\sigma_{i}(\rho^{ss}(\gamma))}{\sigma_{i+1}(\rho^{ss}(\gamma))}\geqslant 2\log|g|_{\Gamma}-R^{\prime}$
@@ 6. Gromov products
$\big{(}g\cdot h)_{\varphi}:=\frac{1}{4}\big{\langle}\varphi,\mu(g)+\mu(g^{-1})+\mu(h)+\mu(h^{-1})-\mu(g^{-1}h)-\mu(h^{-1}g)\big{\rangle}.$
@@
where $\ell=[k_{\ell}e_{1}]$, $V=[k_{V}e_{d}^{\perp}]$ and $k_{V},k_{\ell}\in\mathsf{O}(d)$. The following proposition relates the Gromov product with the limit maps of a representation $\rho$ and will be used in the following sections.
@@
We may write $\rho(\gamma_{n})=w_{n}\exp(\mu(\rho(\gamma_{n})))w_{n}^{\prime}$ and $\rho(\delta_{n})=k_{n}\exp(\mu(\rho(\delta_{n})))k_{n}^{\prime}$ where $w_{n},w_{n}^{\prime},$ $k_{n},k_{n}^{\prime}\in\mathsf{PO}(d)$. Since $\rho$ is $P_{1}$-divergent, $\lim_{n}\frac{\sigma_{d}(\rho(\gamma_{n}))}{\sigma_{j}(\rho(\gamma_{n})}=\lim_{n}\frac{\sigma_{d}(\rho(\delta_{n}))}{\sigma_{j}(\rho(\delta_{n})}=0$ for $11\leq j\leq d-$ 1. Recall that $E_{ij}$ denotes the $d\times d$ elementary matrix with $1$ on the $(i,j)$-entry. Then we notice that
@@
$=\lim_{n\rightarrow\infty}\Bigg{(}\Bigg{|}\Bigg{|}\Big{(}(k_{n}^{\prime})^{-1}\textup{diag}\Big{(}\frac{\sigma_{d}(\rho(\delta_{n}))}{\sigma_{1}(\rho(\delta_{n}))},\ldots,1\Big{)}k_{n}^{-1}w_{n}\textup{diag}\Big{(}1,\ldots,\frac{\sigma_{d}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))}\Big{)}w_{n}^{\prime}\Bigg{|}\Bigg{|}\cdot$
$\Bigg{|}\Bigg{|}\Big{(}(w_{{n}}^{\prime})^{-1}\textup{diag}\Big{(}\frac{\sigma_{d}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))},\ldots,1\Big{)}w_{{n}}^{-1}k_{{n}}\textup{diag}\Big{(}1,\ldots,\frac{\sigma_{d}(\rho(\delta_{n}))}{\sigma_{1}(\rho(\delta_{n}))}\Big{)}k_{{n}}^{\prime}\Bigg{|}\Bigg{|}\Bigg{)}$
$=\lim_{n\rightarrow\infty}\big{|}\big{|}E_{1d}w_{n}^{-1}k_{n}E_{11}\big{|}\big{|}\cdot\big{|}\big{|}E_{1d}k_{n}^{-1}w_{n}E_{11}\big{|}\big{|}$
@@
Fix $\alpha\in\theta$. By [22, Lem. 3.2], there exists $N_{\alpha}>0$ and an irreducible $\theta$-proximal representation $\tau_{\alpha}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ whose highest weight is $N_{\alpha}\omega_{\alpha}$, $N_{\alpha}\in\mathbb{N}$. Since $\rho$ is $P_{\{\alpha\}}$-Anosov, the representation $\tau_{\alpha}\circ\rho$ is $P_{1}$-Anosov. Note that there exists $C_{1}>0$, depending only on $\tau$, such that
$\Big{|}\log\sigma_{1}(\tau_{\alpha}(g))-N_{\alpha}\big{\langle}\omega_{\alpha},\mu(g)\big{\rangle}\Big{|}\leq C_{1}$
@@
for every $\gamma,\delta\in\Gamma$. Therefore, we may continue by assuming that $\rho$ is semisimple. By using Lemma 2.10, we may further assume that $\tau_{\alpha}\big{(}\rho(\Gamma)\big{)}$ has reductive Zariski closure in $\mathsf{GL}_{d}(\mathbb{R})$ and preserves a properly convex open domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$. Let us fix $x_{0}\in\Omega$. By Lemma 2.8 we can find $C_{4}>0$ such that
$\Big{|}\big{(}\tau_{\alpha}(\rho(\gamma))\cdot\tau_{\alpha}(\rho(\delta))\big{)}_{\varepsilon_{1}}-\big{(}\tau_{\alpha}\big{(}\rho(\gamma)\big{)}x_{0}\cdot\tau_{\alpha}\big{(}\rho(\delta)\big{)}x_{0}\big{)}_{x_{0}}\Big{|}\leqslant C_{4}$ (8)
for every $\gamma,\delta\in\Gamma$. By [17] and [39], since $\tau_{\alpha}\circ\rho$ is $P_{1}$-Anosov, $\tau_{\alpha}\big{(}\rho(\Gamma)\big{)}$ acts cocompactly on a closed convex subset $\mathcal{C}\subset\Omega$. Fix $x_{0}\in\mathcal{C}$. The Svarc–Milnor lemma implies that the orbit map $\gamma\mapsto\tau_{\alpha}(\rho(\gamma))x_{0}$ is a quasi-isometry between the Gromov hyperbolic spaces $(\Gamma,d_{\Gamma})$ and $(\mathcal{C},d_{\Omega})$. In particular, there exist $C_{5},c_{5}>0$ such that for every $\gamma,\delta\in\Gamma$,
@@ 7. Characterizations of Anosov representations
If $\rho$ is $P_{\theta}$-Anosov, the Anosov limit maps of $\rho$ are transverse and dynamics preserving and $\rho$ is $P_{\theta}$-divergent (see Theorem 2.3). Also, the fact that the Anosov limit maps satisfy the Cartan property is contained in [22, Thm. 1.3 (4) & 5.3 (4)].
Now we assume that $\rho$ satisfies (i) and (ii). We first reduce to the case where $\Gamma$ is torsion free. Since $\rho$ is $P_{\theta}$-divergent, every element of the kernel $\textup{ker}(\rho)$ has finite order, hence $\textup{ker}(\rho)$ is finite. The quotient group $H=\Gamma/\textup{ker}(\rho)$ is quasi-isometric to $\Gamma$ and by Selberg’s lemma [36] $H$ contains a torsion free and finite-index subgroup $H_{1}$. It is enough to prove that the induced representation $\hat{\rho}:H_{1}\rightarrow G$ is $P_{\theta}$-Anosov. Notice that $\hat{\rho}$ satisfies the same assumptions as $\rho$ and the source group is torsion free.
@@
which induces the splitting $\sigma_{\ast}\mathcal{E}=\sigma_{\ast}\mathcal{E}^{+}\oplus\sigma_{\ast}\mathcal{E}^{-}$, where ±$\mathcal{E}^{\pm}\subset\mathsf{T}(G/L_{\theta})$ are sub-bundles defined in subsection 2.5. Then we fix $x=[\hat{m}]_{\Gamma}$ and choose an element $h\in G$ so that $\xi^{+}(\tau^{+}(\hat{m}))=hP_{1}^{+}$ and $\xi^{-}(\tau^{-}(\hat{m}))=hP_{1}^{-}$. Let $(t_{n})_{n\in\mathbb{N}}$ be an increasing unbounded sequence and consider a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ such that $(\gamma_{n}\varphi_{t_{n}}(\hat{m}))_{n\in\mathbb{N}}$ lies in a compact subset of $\hat{\Gamma}$. We observe that $\lim_{n}\gamma_{n}^{-1}=\tau^{+}(\hat{m})$ in the bordification $\Gamma\cup\partial_{\infty}\Gamma$. Moreover, observe that we can write $\rho(\gamma_{n}^{-1})=(k_{n}^{\prime})^{-1}w\exp(\mu(\rho(\gamma_{n}^{-1}))wk_{n}^{-1}$, where $w=\sum_{i=1}^{d}E_{i(d+1-i)}\in\mathsf{O}(d)$. Since $\xi^{+}$ is assumed to satisfy the Cartan property and $(\gamma_{n})_{n\in\mathbb{N}}$ is $P_{\theta}$-divergent, up to subsequence, we may assume that $\lim_{n}\Xi_{1}^{+}\big{(}\rho(\gamma_{n}^{-1})\big{)}=\lim_{n}(k_{{n}}^{\prime})^{-1}wP_{\theta}^{+}=hP_{\theta}^{+}$. Equivalently, if $k^{\prime}=\lim_{n}k_{n}^{\prime}$ then $k^{\prime}h=w\begin{pmatrix}[0.7]s&\ast\\ 0&B\end{pmatrix}$ for some $B\in\mathsf{GL}_{d-1}(\mathbb{R})$. Fix $u\in\{0\}\times\mathbb{R}^{d-1}$. Then, since $k^{\prime}(k^{\prime})^{t}=I_{d}$, we observe that
@@
We deduce that $\lim_{n}\frac{||\rho(\gamma_{n})^{\ast}h^{-t}u||}{||\rho(\gamma_{n})^{\ast}h^{-t}e_{1}||}=0$ and hence by Proposition 3.2 (ii) we conclude that
$\lim_{n\rightarrow\infty}\big{|}\big{|}\varphi_{t_{n}}(X_{u}^{-})\big{|}\big{|}_{\varphi_{t_{n}}(x)}=0.$
@@
Assume that conditions (i) and (ii) hold. Let $\tau:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ be an irreducible and $\theta$-proximal representation as in Proposition 2.4. It is enough to prove that $\rho^{\prime}=\tau\circ\rho$ is $P_{1}$-Anosov. By using [22, Thm. 5.3 (1)] (see also Lemma 4.8), there exists a pair of continuous, $\rho^{\prime}$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ satisfying the Cartan property. Let $x,y\in\partial_{\infty}\Gamma$ be two distinct points and $(\gamma_{n})_{n\in\mathbb{N}}$ a sequence of elements of $\Gamma$ with $x=\lim_{n}\gamma_{n}$ and $y=\lim_{n}\gamma_{n}^{-1}$. The second condition, shows that
@@
Conversely, part (i) follows immediately by Theorem 2.3 (i). Note that there is $N\geq 1$ such that $N\omega_{\alpha}$ is the highest weight $\chi_{\tau_{\alpha}}$ of an irreducible proximal representation $\tau_{\alpha}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ such that $N\langle\omega_{\alpha},\mu(h)\rangle=\log\sigma_{1}(\tau_{\alpha}(h))$ for every $h\in G$ (see [22, Lem. 3.2]). By Proposition 1.12 (i) we can find $B,b>0$ such that for every $\alpha\in\theta$ and $\gamma\in\Gamma$ we have
$\big{\langle}\omega_{\alpha},2\mu(\rho(\gamma))-\mu(\rho(\gamma^{2}))\big{\rangle}\leq\big{\langle}\omega_{\alpha},2\mu(\rho(\gamma))+2\mu(\rho(\gamma^{-1}))-\mu(\rho(\gamma^{2}))-\mu(\rho(\gamma^{-2}))\big{\rangle}\leqslant B(\gamma\cdot\gamma^{-1})_{e}+b.$
Therefore, $\big{\langle}\omega_{\alpha},2\mu(g)-\mu(g^{2})\big{\rangle}\geqslant 0$ for every $g\in G$. This concludes the proof of the corollary. ∎

### v4

@@ 5. Property (U), weak eigenvalue gaps and the uniform gap summation property
$\rho(g^{n})=\begin{pmatrix}I_{2}&\big(\sum_{i=0}^{n}A^{i}\big)^{t}B\\ 0&A^{n}\end{pmatrix}$
@@
Similarly as in the previous case, we deduce that $\rho(\textup{ker}(\rho_{V}))$ (and hence $\textup{ker}(\rho_{V})$) is virtually nilpotent and finitely generated. The quotient $\Gamma/\textup{ker}(\rho_{V})$ is abelian, so $\Gamma$ has to be virtually polycyclic. Since $|\partial_{f}\Gamma|>1$, a theorem of Floyd [21, p. 211] implies that $\Gamma$ has two ends, so $\Gamma$ is virtually cyclic. Since $\Gamma$ is assumed not to be virtually nilpotent, this is again a contradiction, hence $\partial_{f}\Gamma$ cannot contain two points.
It follows that $|\partial_{f}\Gamma|\geq 3$. Therefore, Theorem 1.7 shows that $\Gamma$ satisfies strong property (U). ∎
Proof of Theorem 1.6..
Suppose that (i) holds, i.e $\rho$ is $P_{i}$-Anosov. Then (ii) holds since the Floyd boundary identifies with the Gromov boundary of $\Gamma$. Moreover, by Theorem 2.3 and Proposition 2.6, (iii) and (iv) hold true for any semisimplification $\rho^{ss}$ of the $P_{i}$-Anosov representation $\rho$. Now let us prove the other implications. We assume that there exists $\varepsilon>0$ such that for every $\gamma\in\Gamma$,
By [33, Prop. 4.12] it is enough to prove that $\Gamma$ satisfies strong property (U).
$\textup{(ii)}\Rightarrow\textup{(i)}$. We first observe that for every element $g\in\textup{ker}(\rho)$ we have $|g|_{\infty}=0$. We next show that $N:=\textup{ker}\rho$ is finite. If not, $N$ is an infinite normal subgroup of $\Gamma$ and $\Lambda(N)=\partial_{f}\Gamma$ since $\Gamma$ acts minimally on $\partial_{f}\Gamma$. By [32, Thm. 1] there exists a non cyclic free subgroup $H$ of $N$ with $|\Lambda(H)|\geq 3$. In particular, by Theorem 1.7 we can find $\gamma\in H$ such that $|\gamma|_{\infty}>0$. This is a contradiction since $\gamma\in N$. It follows that $N$ is finite.
The Floyd boundary of $\Gamma^{\prime}=\Gamma/N$ is non-trivial since $\Gamma^{\prime}$ is quasi-isometric to $\Gamma$. Note that the representation $\rho$ induces a faithful representation $\rho^{\prime}:\Gamma^{\prime}\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ which also has a weak uniform $i$-gap in eigenvalues. Selberg’s lemma [40] implies that $\Gamma^{\prime}$ is virtually torsion-free, thus, by Theorem 1.7, $\Gamma^{\prime}$ satisfies strong property (U). We conclude that $\Gamma^{\prime}$ and $\Gamma$ are word hyperbolic and $\rho$ is $P_{i}$-Anosov.
$\textup{(iii)}\Rightarrow\textup{(i)}$. If $\Gamma$ is virtually nilpotent, Lemma 5.2 implies that $\Gamma$ is virtually cyclic, contradicting our assumption. Since $\Gamma$ is not virtually nilpotent and $\rho_{1}$ satisfies the uniform gap summation property, $\Gamma$ has to satisfy strong property (U) by Corollary 1.8. Therefore, (i) holds.
$\textup{(iv)}\Rightarrow\textup{(i)}$. Let $\rho^{ss}$ be a semisimplification of $\rho$. By Proposition 2.6, $\lambda(\rho(g))=\lambda(\rho^{ss}(g))$ for every $g\in\Gamma$, hence there exists $c_{2}>0$, depending only on $\rho_{2}$, such that
[…]
for every $\gamma\in\Gamma$. By Theorem 2.5 there exists a finite subset $F$ of $\Gamma$ and $C>0$ such that for every $\gamma\in\Gamma$ there is $w\in F$ with
[…]
In particular, we may choose $R>0$ such that for every $\gamma\in\Gamma$,
[…]
By assumption, for all but finitely many $\gamma\in\Gamma$ we have
[…]
so there exists $R^{\prime}>0$ such that
$\log\frac{\sigma_{i}(\rho^{ss}(\gamma))}{\sigma_{i+1}(\rho^{ss}(\gamma))}\geqslant 2\log|\gamma|_{\Gamma}-R^{\prime}$
for all $\gamma\in\Gamma$ non-trivial. In particular, the semisimplification $\rho^{ss}$ of $\rho$ satisfies the uniform gap summation property. Therefore, since $\rho^{ss}$ has a weak uniform $i$-gap in eigenvalues, by implication $\textup{(iii)}\Rightarrow\textup{(i)}$, $\rho^{ss}$ is $P_{i}$-Anosov and $\Gamma$ is word hyperbolic. In particular, $\rho$ is $P_{i}$-Anosov. ∎
@@ 6. Gromov products
$\big(g\cdot h)_{\varphi}:=\frac{1}{4}\varphi\Big(\mu(g)+\mu(g^{-1})+\mu(h)+\mu(h^{-1})-\mu(g^{-1}h)-\mu(h^{-1}g)\Big).$
@@
where $\ell=[k_{\ell}e_{1}]$, $V=[k_{V}e_{d}^{\perp}]$, $k_{V},k_{\ell}\in\mathsf{O}(d)$ and $\langle\cdot,\cdot\rangle$ is the standard inner product. Recall that a representation $\rho:\Gamma\rightarrow\mathsf{PGL}_{d}(\mathbb{R})$ is called $P_{1}$-divergent if $\lim\frac{\sigma_{1}(\rho(\gamma_{n}))}{\sigma_{2}(\rho(\gamma_{n}))}=\infty$ as $|\gamma|_{\Gamma}\rightarrow\infty$.
The following proposition relates the Gromov product with the limit maps of a representation $\rho$ and will be used in the following sections.
@@
We may write $\rho(\gamma_{n})=w_{n}\exp(\mu(\rho(\gamma_{n})))w_{n}^{\prime}$ and $\rho(\delta_{n})=k_{n}\exp(\mu(\rho(\delta_{n})))k_{n}^{\prime}$ where $w_{n},w_{n}^{\prime},$ $k_{n},k_{n}^{\prime}\in\mathsf{PO}(d)$. Since $\rho$ is $P_{1}$-divergent, $\lim_{n}\frac{\sigma_{d}(\rho(\gamma_{n}))}{\sigma_{j}(\rho(\gamma_{n}))}=\lim_{n}\frac{\sigma_{d}(\rho(\delta_{n}))}{\sigma_{j}(\rho(\delta_{n}))}=0$ for $1\leq j\leq d-1$. Recall that $E_{ij}$ denotes the $d\times d$ elementary matrix with $1$ on the $(i,j)$-entry. Then we notice that
@@
$=\lim_{n\rightarrow\infty}\Bigg(\Bigg|\Bigg|(k_{n}^{\prime})^{-1}\textup{diag}\Bigg(\frac{\sigma_{d}(\rho(\delta_{n}))}{\sigma_{1}(\rho(\delta_{n}))},\ldots,1\Bigg)k_{n}^{-1}w_{n}\textup{diag}\Bigg(1,\ldots,\frac{\sigma_{d}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))}\Bigg)w_{n}^{\prime}\Bigg|\Bigg|\cdot$
$\Bigg|\Bigg|(w_{{n}}^{\prime})^{-1}\textup{diag}\Bigg(\frac{\sigma_{d}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))},\ldots,1\Bigg)w_{{n}}^{-1}k_{{n}}\textup{diag}\Bigg(1,\ldots,\frac{\sigma_{d}(\rho(\delta_{n}))}{\sigma_{1}(\rho(\delta_{n}))}\Bigg)k_{{n}}^{\prime}\Bigg|\Bigg|\Bigg)$
[…]
@@
Fix $\alpha\in\theta$. By [42, Thm. 7.2], there exists $N_{\alpha}>0$ and an irreducible $\theta$-proximal representation $\tau_{\alpha}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ whose highest weight is $N_{\alpha}\omega_{\alpha}$, $N_{\alpha}\in\mathbb{N}$. Since $\rho$ is $P_{\{\alpha\}}$-Anosov, the representation $\tau_{\alpha}\circ\rho$ is $P_{1}$-Anosov. There exists $C_{1}>0$, depending only on $\tau_{\alpha}$, such that
$\Big|\log\sigma_{1}(\tau_{\alpha}(g))-N_{\alpha}\omega_{\alpha}(\mu(g))\Big|\leq C_{1}$
@@
for every $\gamma,\delta\in\Gamma$. Therefore, we may continue by assuming that $\rho$ is semisimple. By using Lemma 2.10, we may further assume that $\tau_{\alpha}\big(\rho(\Gamma)\big)$ has reductive Zariski closure in $\mathsf{GL}_{d}(\mathbb{R})$ and preserves a properly convex open domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$. Let us fix $x_{0}\in\Omega$. By Lemma 2.8 we can find $C_{4}>0$ such that for every $\gamma,\delta\in\Gamma$,
$\Big|\big(\tau_{\alpha}(\rho(\gamma))\cdot\tau_{\alpha}(\rho(\delta))\big)_{\varepsilon_{1}}-\big(\tau_{\alpha}\big(\rho(\gamma)\big)x_{0}\cdot\tau_{\alpha}\big(\rho(\delta)\big)x_{0}\big)_{x_{0}}\Big|\leqslant C_{4}.$ (10)
By [19] and [44], since $\tau_{\alpha}\circ\rho$ is $P_{1}$-Anosov, $\tau_{\alpha}\big(\rho(\Gamma)\big)$ acts cocompactly on a closed convex subset $\mathcal{C}\subset\Omega$. Fix $x_{0}\in\mathcal{C}$. The Svarc–Milnor lemma implies that the orbit map $\gamma\mapsto\tau_{\alpha}(\rho(\gamma))x_{0}$ is a quasi-isometry between the Gromov hyperbolic spaces $(\Gamma,d_{\Gamma})$ and $(\mathcal{C},d_{\Omega})$. In particular, there exist $C_{5},c_{5}>0$ such that for every $\gamma,\delta\in\Gamma$,
@@ 7. Characterizations of Anosov representations
If $\rho$ is $P_{\theta}$-Anosov, the Anosov limit maps of $\rho$ are transverse and dynamics preserving and $\rho$ is $P_{\theta}$-divergent (see Theorem 2.3). Also, the fact that the Anosov limit maps satisfy the Cartan property is contained in [24, Thm. 1.3 (4) & 5.3 (4)].
Now we assume that $\rho$ satisfies (i) and (ii). We first reduce to the case where $\Gamma$ is torsion-free. Since $\rho$ is $P_{\theta}$-divergent, every element of the kernel $\textup{ker}(\rho)$ has finite order, hence $\textup{ker}(\rho)$ is finite. The quotient group $\Gamma_{1}=\Gamma/\textup{ker}(\rho)$ is quasi-isometric to $\Gamma$ and by Selberg’s lemma [40] $\Gamma_{1}$ contains a torsion-free and finite-index subgroup $\Gamma_{2}$. It is enough to prove that the induced representation $\hat{\rho}:\Gamma_{2}\rightarrow G$ is $P_{\theta}$-Anosov. Notice that $\hat{\rho}$ satisfies the same assumptions as $\rho$ and the source group is torsion-free.
@@
[…]
@@
[…]
@@
Assume that conditions (i) and (ii) hold. Let $\tau_{\theta}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ be an irreducible and $\theta$-proximal representation as in Proposition 2.4. In order to show that $\rho$ is $\theta$-Anosov, it suffices to check that $\rho^{\prime}=\tau_{\theta}\circ\rho$ is $P_{1}$-Anosov. By using [24, Thm. 5.3 (1)] (see also Lemma 4.8), there exists a pair of continuous, $\rho^{\prime}$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ satisfying the Cartan property. Let $x,y\in\partial_{\infty}\Gamma$ be two distinct points and $(\gamma_{n})_{n\in\mathbb{N}}$ a sequence of elements of $\Gamma$ with $x=\lim_{n}\gamma_{n}$ and $y=\lim_{n}\gamma_{n}^{-1}$. Condition (ii), shows that
@@
Conversely, part (i) follows immediately by Theorem 2.3 (i). Note that there is $N_{\alpha}\geq 1$ such that $N_{\alpha}\omega_{\alpha}$ is the highest weight of an irreducible proximal representation $\tau_{\alpha}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$. There is a constant $C_{0}>0$, depending only on $\tau_{\alpha}$ such that $\big|N_{\alpha}\omega_{\alpha}(\mu(h))-\log\sigma_{1}(\tau_{\alpha}(h))\big|\leq C_{0}$ for every $h\in G$. By Proposition 1.12 (i) and using the fact that for every $h\in G$, $\omega_{\alpha}(2N_{\alpha}\mu(h)-N_{\alpha}\mu(h^{2}))\geq 2\log\sigma_{1}(\tau_{\alpha}(h))-\log\sigma_{1}(\tau_{\alpha}(h^{2}))-3C_{0}\geq-3C_{0}$, we can find $B,b>0$ such that for every $\alpha\in\theta$ and $\gamma\in\Gamma$ we have
$\omega_{\alpha}\big(2\mu(\rho(\gamma))-\mu(\rho(\gamma^{2}))\big)$ $\leq 3C_{0}N_{\alpha}^{-1}+\omega_{\alpha}\big(2\mu(\rho(\gamma))+2\mu(\rho(\gamma^{-1}))-\mu(\rho(\gamma^{2}))-\mu(\rho(\gamma^{-2}))\big)$
$\leqslant B(\gamma\cdot\gamma^{-1})_{e}+b.$
This concludes the proof of the corollary. ∎